OpenStudy (anonymous):

Since the opening night, attendance at Play A has increased steadily, while attendance at Play B first rose and then fell. Equations modeling the daily attendance y at each play are shown below, where x is the number of days since the opening night. On what day(S) was the attendance the same at both plays? what was the attendance? Play A: y = 16x + 150 Play B: y = -x^2 60x - 10 answer choices: The attendance was the same on day 40. The attendance was 790 at both plays that day. The attendance was the same on day 4. The attendance was 214 at both plays that day. The attendance w

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

set the y's equal and solve

OpenStudy (anonymous):

okay

OpenStudy (triciaal):

this will give the day the attendance the same is that a "+ 60x"

OpenStudy (triciaal):

16x + 150 = -x^2 + 60x -10 x^2 -60x +16x + 150 + 10 = 0

OpenStudy (anonymous):

so the answer is: the attendance was never the same on both plays?

OpenStudy (anonymous):

can I ask another question @triciaal

OpenStudy (triciaal):

sure but before you go to the next ? how did you decide this one

OpenStudy (anonymous):

Because of the step which shows it was 0, that meant that the answer would be none, and they never did. that was the last answer choice, it just did fully come out.

OpenStudy (triciaal):

did you plug the choices and check you did not verify +60x

OpenStudy (anonymous):

so your saying I'm wrong?

OpenStudy (triciaal):

i am saying did not find x and therefore did not actually check to confirm did not plug the given numbers to check

OpenStudy (anonymous):

Wait so you don't know if I'm right? I'm confused now, and I don't usually go back and check like that how would I do that?

OpenStudy (triciaal):

I set the y's equal to find the value of x but the eqn written was not clarified so i did not solve for x; you did not answer

OpenStudy (triciaal):

what is the next question?

OpenStudy (anonymous):

oh so was I right? or not?

OpenStudy (anonymous):

let me write the next question

OpenStudy (triciaal):

the answer is b it I use +60x

OpenStudy (triciaal):

look at the question did you miss a plus sign that's what i need to know and you did not answer

OpenStudy (anonymous):

yes on Play B

OpenStudy (anonymous):

there was a missing + sign ur right

OpenStudy (anonymous):

14. What ate tje solutions to the system? y = x^2 - 6x + 7 y = -x + 13 (-1, 14) and (6, 14) (-1, 7) and (6, 14) (-1, 13) and (6, 7) no solution.

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

set the y's equal

OpenStudy (triciaal):

x^2 -6x + 7 = -x + 13

OpenStudy (triciaal):

add x -13 to each side

OpenStudy (triciaal):

group like terms

OpenStudy (triciaal):

x^2 -5x -6 = 0

OpenStudy (triciaal):

what are the factors of -6 that add to -5

OpenStudy (triciaal):

lets use -6 and + 1 x^2 - 6x + x -6 = 0 x(x + 1)-6(x+1) = 0 (x+1)(x-6)=0

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

okay

OpenStudy (triciaal):

x= -1, x = 6 you have the x value put in the 2nd eqn and find y

OpenStudy (triciaal):

i dont see the ans I got (-1, 14) and (6, 7)

OpenStudy (anonymous):

its there, its the first option

OpenStudy (triciaal):

your first option has (6, 14)

OpenStudy (anonymous):

oops haha, on my test it says: (-1, 14) and (6, 7) I copied wrong lol its supposed to be the first choice.

OpenStudy (anonymous):

can we do another?

OpenStudy (triciaal):

thank you

OpenStudy (triciaal):

last one for now i have to get going

OpenStudy (anonymous):

haha no prob

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

15. Simplify the radical expression. \[\sqrt{72} \] \[2\sqrt{6}\] \[6\sqrt{2}\] \[-6\sqrt{2}\] 12

OpenStudy (anonymous):

alright there it is and I will be back in just a moment

OpenStudy (triciaal):

what are the prime factors of 72

OpenStudy (triciaal):

rt 72 = rt of 2*2*2*3*3 2*3 *rt2 = 6rt2

OpenStudy (triciaal):

remember rt2 *rt2 = 2 and rt3*rt3 = 3

OpenStudy (triciaal):

with practice you will learn to factor out perfect sq 72 = 4 * 2 * 9 72 = 2*36 perfect sqs 4,9, 36

OpenStudy (triciaal):

choice c nonetheless later

OpenStudy (anonymous):

thank you!! have a nice day sorry i had to leave for a moment.